OpenStudy (mtalhahassan2):

Need help!! For each of the following values of sin theta determine the measure of theta if pie/2

OpenStudy (mtalhahassan2):

what you mean my the choices

OpenStudy (anonymous):

\[\frac{\sqrt{2}}{2} = \sin( \theta)\]

OpenStudy (anonymous):

is this what you would like to know?

OpenStudy (mtalhahassan2):

yes

OpenStudy (anonymous):

if you have a calculator you can check it by taking the inverse sin of root 2 over 2 otherwise by unit circle defintion we know that it is \[\frac{\pi}{4}\]

OpenStudy (anonymous):

that is not in our domain so we have to find one in the second quadrant

OpenStudy (anonymous):



OpenStudy (anonymous):

so there is two angles that would work in the domain

OpenStudy (anonymous):



OpenStudy (anonymous):

whats halfway between pi/2 and pi

OpenStudy (mtalhahassan2):

3pie/2

OpenStudy (anonymous):

no that is at the bottom

OpenStudy (anonymous):

maybe this will help. its between \[\frac{2\pi}{4} and \frac{4\pi}{4}\]

OpenStudy (mtalhahassan2):

3pie/4

OpenStudy (anonymous):

yes

OpenStudy (mtalhahassan2):

I have a question like how you find the pie/4

OpenStudy (mtalhahassan2):

so basically the angle be in the quardant 2 and 3

OpenStudy (mtalhahassan2):

right

OpenStudy (mtalhahassan2):

because the given info tell us that it be between 90 to 270